1) Random variable x has a uniform distribution defined by the probability density function below. Determine the probability that x has a value of at least 220.

    f(x) =           1/100 for values of x between 200 and 300, and

                       0   everywhere else

A) 0.65

B) 0.80

C) 0.75

D) 0.60

2) The method of sampling that ensures that every subgroup of interest in a particular study is represented in the sample is called:

A) systematic random sampling

B) stratified random sampling

C) grouped random sampling

D) cluster random sampling

3) You want to estimate the average GMAT score for all students who took the Logitano GMAT Preparation course during the past 2 years. You select a random sample of 100 such students from a comprehensive list of all students who took the course over the last two years and find that the average GMAT score for the sample was 1040. Suppose now you build a 95% confidence interval estimate of the average score for the population of Logitano students represented here and the interval turns out to be 1040 ± 10 points. How would you interpret your result?

A) We can be 95% confident that the interval 1030 to 1050 contains the average score for the entire population of Logitano students.

B) The score for 95% of all Logitano students will be between 1030 and 1050.

C) There is a 5% chance that not all Logitano students scored between 1030 and 1050.

D) We can be 95% confident that all of Logitano students scored between 1030 and 1050.

4) The sampling distribution of the sample proportion can be treated as a normal distribution when which of the following conditions holds?

A) n(pi) > or = 5

B) n(1 – pi) > or = 5

C) pi has a uniform distribution

D) both a and b
Answer #1

1 )P(X>220) =1-(220-200)/100 =1-0.2 =0.80

option B

2) B) stratified random sampling

3)

A) We can be 95% confident that the interval 1030 to 1050 contains the average score for the entire population of Logitano students.

4) D) both a and b